OpenStudy (anonymous):

one of the vertices of the hyperbola (x-3)^2/11-(y-5)^2/11=1

OpenStudy (jdoe0001):

$$ \text{distance from the center of the hyperbola}\\ \text{ to any of the foci is}\\ \sqrt{a^2+b^2} $$ so, get that value and add it to, in this case to the "h" from (h,k) center coordinates

OpenStudy (anonymous):

well a=2 and b= sqrt of 2 so how would I plug that in

OpenStudy (jdoe0001):

btw, I'm guessing you meant to say \(11^2 \) for denominator, no "11" :|

OpenStudy (anonymous):

no just 11

OpenStudy (jdoe0001):

hmm, ok, well, then just \(\sqrt{11^2+11^2} \), what does that give you?

OpenStudy (anonymous):

sqrt 242 or 15.5

OpenStudy (anonymous):

but I don't think that's righhht. Its not in my answers. Im looking for one of the vertices of the hyperbola

OpenStudy (jdoe0001):

ok, what's your center? your (h,k)

OpenStudy (anonymous):

center is 1,3

OpenStudy (jdoe0001):

1,3?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

hehe, do you know what your (h,k) values are from the equation?

OpenStudy (jdoe0001):

$$ \cfrac{(x-\color{red}{3})^2}{11}-\cfrac{(y-\color{red}{5})^2}{11}=1 $$

OpenStudy (anonymous):

yes perfect

OpenStudy (jdoe0001):

well, since the rational that is POSITIVE is the one with the "x" in the numerator, that means the hyperbola is moving HORIZONTALLY

OpenStudy (jdoe0001):

so, your focus would be at \(\huge \color{red}{h}\pm 15.5 \)

OpenStudy (anonymous):

yes except remember that (y-5)^2/7 not 11

OpenStudy (jdoe0001):

foci for plural rather :)

OpenStudy (anonymous):

well here are my options: a (5-sqrt11,3) b.(5-sqrt11,-3) (3-sqrt11,-5) d.(3+sqrt11,5)

OpenStudy (jdoe0001):

well... you want the vertices.... hmmm ok

OpenStudy (jdoe0001):

shoot.... that's even simpler.... from the center, which you already have, the distance is THE DENOMINATOR UNDER THE POSITIVE FRACTION :)

OpenStudy (anonymous):

Im confused could you show me

OpenStudy (jdoe0001):

the value not squared :), so \(\sqrt{11}\)

OpenStudy (jdoe0001):

$$ \cfrac{(x-\color{red}{3})^2}{(\color{blue}{\sqrt{11}})^2}-\cfrac{(y-\color{red}{5})^2}{(\sqrt{11})^2}=1 $$

OpenStudy (jdoe0001):

that's pretty much the hyperbola equation in standard form, so you want the "a" element, or the blue one in this case

OpenStudy (anonymous):

yess so would it 3-sqrt 11 or 5-squrt 11

OpenStudy (jdoe0001):

well, of the 2 fractions in the equation, which one is the POSITIVE one? the one with the "x" in it or the one with the "y" in it?

OpenStudy (anonymous):

the one with the x

OpenStudy (jdoe0001):

that means the hyperbola is moving HORIZONTALLY or opening over the "x", and the vertices will be there

OpenStudy (jdoe0001):

so for (h,k) or (x,y) you have (3,5) for your center, so \(h\pm 15.5\)

OpenStudy (jdoe0001):

wait... it wasn't 15.5.... ahemm, 3.3 rather

OpenStudy (jdoe0001):

or \(\sqrt{11}\)

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

so the answerrr would be A

OpenStudy (jdoe0001):

is it?, what's your "h" value again?

OpenStudy (jdoe0001):

from the center form (h,k) or what is your "x" coordinate for the center for the hyperbola?

OpenStudy (jdoe0001):